### v3

# Topological Defects and Geometrical Frustrations in Fourier Photonic Simulator

Abstract
The XY models with continuous spin orientation play a pivotal role in understanding topological phase transitions and emergent frustration phenomena, such as superconducting and superfluid phase transitions. However, the complex energy landscapes arising from frustrated lattice geometries and competing spin interactions make these models computationally intractable. To address this challenge, we designed a programmable photonic spin simulator capable of emulating the XY models with tunable lattice geometries and spin couplings, allowing systematic exploration of their statistical properties. We experimentally observed the Berezinskii-Kosterlitz-Thouless transition in a square-lattice XY model with nearest-neighbor interactions and determined its critical temperature. Expanding to frustrated systems, we implemented the approach in triangular and honeycomb lattices, uncovering sophisticated phase transitions and frustration effects, which were consistent with theoretical predictions. This versatile platform opens avenues for probing unexplored XY model phenomena across diverse geometries and interaction schemes, with potential applications in solving complex optimization and machine learning problems.
The XY models [1] provide profound insight into complex dynamics in multidisciplines in condensed matter physics [2, 3], machine learning [4, 5], and continuous optimization [6]. These models are characterized by continuous spin orientations in two-dimensional lattices, which host a topological phase transition at finite temperature, that is, the Berezinskii-Kosterlitz-Thouless (BKT) phase transition [7]. Recently, it has been successfully applied to explain phase transitions in superconducting [2] and superfluid [3] films, which are more consistent with the experimental results than the Bardeen–Cooper–Schrieffer (BCS) theory. To explore physical phenomena from XY models, sampling spin configurations from the phase space according to the Boltzmann distribution is required [8, 9]. This sampling process is computationally daunting due to the exponential increase in the number of spin configurations with the number of spins [10]. In addition, the energy landscape of the XY models is rugged arising from frustrated lattice geometries and various spin interactions, resulting in the extremely long relaxation of sampling procedure for generating effective spin configurations [11]. As a result, the investigation of XY models is resource intensive and challenging to expand to a large scale on computers.
To mitigate this difficulty, analog simulators have emerged as powerful alternatives to replace conventional electronic computers. Platforms such as optical parametric oscillators [11, 12], superconducting quantum circuits [13], spatial light modulators (SLMs) [14, 15, 16, 17, 18, 19, 20], degenerate cavities [21, 22], nanolasers [23] and polariton condensates [24, 25] are used to encode spins, allowing efficient sampling and optimization. Among these, the free-space light-based photonic simulators stand out for their accessibility, scalability, and compatibility to existing algorithms [14, 16, 15], owing to the use of readily segmentable and accessible wavefront for encoding spins. Through this simulator, the phase transition of spin glass models can be experimentally observed by annealing samplings[17, 18, 19], and combinatorial optimization problems can be solved by searching the ground states of spin glass models [14, 15].
Despite its versatility, previous studies mostly focused on the simulation of Ising models with binary value spins. Recently, continuous spins on square lattice without frustration have also been investigated [26, 27]. However, the encoding and sampling of spin configurations in XY models with programmable lattice geometries and spin interactions are still challenging. These continuous spin models are more general in thermodynamics and can be directly applied to practical continuous optimization problems [4, 5], for instance, continuous graph coloring [28], angle synchronization problems [29] and max $k$-cut [30]. Toward this goal, a photonic simulator has been constructed with delicate but sophisticated design and algorithm [20], where the BKT transition was studied. However, its scalability in lattice size is limited and the generalizability to different lattice geometries has not been substantiated.
Figure 1: The principle of FPS. The XY spins $\bm{S}_{j}$ with angle $\theta_{j}$ are encoded to the wavefront phase $\varphi_{j}$ of a laser by an SLM. The modulated light passes through a lens, which performs an optical Fourier transform, and is finally detected by a camera on the back focal plane of the lens. A pointwise multiplication (Mult.) or Hadamard product is implemented between the light intensity $I$ and the Fourier mask $I_{M}$, which is equivalent to a convolution (Conv.) operation in the real space. Finally, the Hamiltonian of the XY model is obtained from the summation over all values of the products. Here, the Fourier mask has the effect of a filter for spin interactions, which can select the Hamiltonian of arbitrary range spin interactions. The MCMC algorithm is introduced to update the spin configurations by using the optically obtained Hamiltonian, thereby sampling spin configurations that obey the Boltzmann distribution. These samples can be employed to obtain important quantities of the XY models, such as the correlation function, helicity modulus and specific heat.
In this Letter, we propose and implement a scalable and reconfigurable Fourier photonic simulator (FPS) to experimentally investigate the BKT phase transitions and study the frustration of XY models with programmable lattice geometries and spin interactions. By using an SLM that can simultaneously modulate the phase and amplitude of light, we encode spins into the phase of a light wavefront and engineer spin interactions via a reconfigurable Fourier mask [16]. To validate FPS, we constructed a ferromagnetic XY model on a square lattice, successfully observed the BKT phase transition, and determined the transition temperature. We further extend the approach to frustrated systems, including triangular and honeycomb lattices with competing antiferromagnetic spin interactions [31]. The triangular lattice exhibits a dual BKT and Ising transition, while the honeycomb lattice hosts various ground states, depending on the interaction parameters. These findings highlight the interplay between lattice geometry, frustration, and topology and demonstrate the versatility of FPS.
Fourier-mask enabled photonic spin simulator–The Hamiltonian of XY models is
$\mathcal{H}$ $=-\sum_{(j,\,h)}J_{jh}\bm{S}_{j}\cdot\bm{S}_{h}$
$=-\sum_{\alpha\in\mathbb{Z}^{+}}J_{\alpha}\sum_{(j,\,h)_{\alpha}}\xi_{j}\xi_{h}\cos\left(\theta_{j}-\theta_{h}\right)~,$ (1)
where $\bm{S}=\left(\cos\theta,~\sin\theta\right)$ with $\theta\in\left[0,~2\pi\right)$ is the spin of XY models [32], and $J_{jh}$ denotes the interaction strength between the $j$-th and $h$-th spins. $J_{\alpha}$ is the interaction strength of the $\alpha$ range between neighbors (with $\alpha=1$ being the nearest neighbors). $\xi\in[-1,1]$ is a random variable which characterizes the disorder strength in Mattis-type spin glass models [33]. The $(j,h)_{\alpha}$ under the summation designates spin pairs, such as $(j,h)_{1}=\left\langle j,h\right\rangle$ for the nearest-neighbor (NN) spin pairs and $(j,h)_{2}=\left\langle\left\langle j,h\right\rangle\right\rangle$ for the next-nearest-neighbor (NNN) spin pairs.
In the FPS, the SLM pixel arrays correspond to the spin lattices with the spin orientation angles $\theta$ encoded in the modulated phases of light $\varphi$, and the disorder strength $\xi$ encoded in the light field amplitude $A$ (see FIG. 1 and End Matter). The light reflected from the SLM is transformed by a lens to its Fourier plane, where we place a camera to measure the light intensity $I(\bm{u})$, with $\bm{u}$ being the coordinates on the camera. By specifically designing a mask $I_{M}(\bm{u})$ and make integration of $I(\bm{u})I_{M}(\bm{u})$ we can obtain the Hamiltonian of the spins with various XY interactions in Eq. (1) (see End Matter and the Supplementary Materials [34] for more details).
Figure 2: Observing the BKT transition of ferromagnetic XY model on the square lattice. (a) The lattice structure of a square lattice, the spin arrangement on the SLM, and the Fourier mask for NN interactions with periodic boundary conditions, respectively. $a$ is the lattice constant. The blue arrows represent spin coupling. (b) The helicity modulus ($Y$) as functions of temperature across different lattice sizes ($L\times L$$\times8$, $\times10$, $\times16$, $\times20$). The solid gray line represents $\tilde{Y}=\frac{2}{\pi}T$. The BKT transition is identified by the intersection of $Y(T)$ and $\tilde{Y}(T)$. The dashed gray line indicates $T\approx 0.8922$, which is the $T_{\text{BKT}}$ of infinite lattice. Insert provides a zoomed-in view near the BKT transition temperature $T_{\text{BKT}}$. (c) The correlation function $G(r)$ with respect to the distance $r$ between spins at $T=1.2>T_{\text{BKT}}$ (top) and $T=0.8<T_{\text{BKT}}$ (bottom). Above $T_{\text{BKT}}$, $G(r)$ fits the exponential decay model better with $R^{2}=0.9926$, while the power-law decay mode fits better below $T_{\text{BKT}}$ with $R^{2}=0.9780$.
This strategy is compatible with the free arrangement of spins on the SLM, enabling special lattice configurations beyond square lattices. For instance, honeycomb and kagome lattices can be constructed by introducing vacancies and displacements of spins on the SLM. Since the SLM we used in the experiment can also modulate light intensity, we can introduce vacancy sites by setting the light intensity on those pixels to be zero. After obtaining the Hamiltonian, we use the Markov chain Monte Carlo (MCMC) sampling algorithm, e.g., Metropolis-Hastings [44] and parallel tempering [45], to update the spin configurations and to study XY models across a range of lattice geometries. The sampling loop of FPS is shown in FIG. 1 (see [34] for algorithm details).
Ferromagnetic XY model on the square lattice–We first construct the ferromagnetic XY model on the square lattice with NN interaction, as shown in FIG. 2(a). In this case, the $J_{\alpha}$ in Eq. (1) is $J\delta(\alpha-1)$ where $J$ is the interaction strength constant. Unlike the Ising model, the XY model has no phase transition from spontaneous symmetry breaking [46, 47], but possesses a BKT phase transition driven by interactions between topological defects [48]. These topological defects featured by vortices have significantly different dynamics above and below a phase transition temperature $T_{\text{BKT}}/J\approx 0.8922$ [49, 7]. Vortices with opposite charges, whose spins rotate clockwise and anti-clockwise, respectively, are tightly bound at low temperatures and dissociate above $T_{\text{BKT}}$ [43].
To study BKT phase transition and determine $T_{\text{BKT}}$, we introduce the helicity modulus, $Y(T)=\frac{1}{N}\left[\frac{1}{2}\langle\mathcal{H}\rangle-\beta\left(\langle I_{x}^{2}\rangle-\langle I_{x}\rangle^{2}\right)\right]$, where $\beta=(k_{B}T)^{-1}$ is the inverse temperature with the Boltzmann constant $k_{B}$ (we set $k_{B}=1$ for simplicity), and $N$ is the number of spins. $I_{x}=-\sum_{\langle j,\,h\rangle_{x}}\sin(\theta_{h}-\theta_{j})$ is the spin current along the $x$ direction, and $\langle\cdots\rangle_{x}$ denotes the NN spin pairs along the $x$ direction, $\langle I_{x}^{2}\rangle-\langle I_{x}\rangle^{2}$ is fluctuation of the spin current. Predicted by the renormalization group theory [7], the value of the helicity modulus with an infinite lattice size has a jump from zero to $\frac{2}{\pi}T_{\text{BKT}}$ at the BKT phase transition point. Consequently, the intersection of $Y(T)$ and $\tilde{Y}(T)=\frac{2}{\pi}T$ can be used to determine the value of $T_{\text{BKT}}$. Especially, according to the scaling theory for finite-size systems, the BKT transition temperature increases when we decrease the lattice size [50]. To verify this phenomenon, we simulate XY models on square lattices of various sizes ($N=8\times 8$, $10\times 10$, $16\times 16$, and $20\times 20$) by FPS. As shown in FIG. 2(b), the intersections converge to $T/J=0.8922$ from above when we increase the lattice size, consistent with the theoretical prediction.
The BKT phase transition can be characterized from another perspective. The correlation function $G(\bm{r})=\langle\bm{S}(\bm{x}_{j})\cdot\bm{S}(\bm{x}_{j}+\bm{r})\rangle$ scales differently above and below the $T_{\text{BKT}}$ [51]. When $T>T_{\text{BKT}}$, $G(\bm{r})$ decays exponentially because the vortices are randomly distributed. When $T<T_{\text{BKT}}$, vortex pairs with opposite charges are tightly bound, such that the $G(\bm{r})$ decays in a power-law manner, indicating a quasi-long range order. In FIG. 2(c), we show $G(\bm{r})$ for the two different topological phases at $T/J=1.2$ and $0.8$. The exponential decay function fits better for the high-temperature phase, while the power-law decay function fits better for the low-temperature phase. We also calculated the number of free vortices as a function of the temperature, as well as the change of the topological invariants (see [34]), which are coherent with theoretical predictions. Therefore, FPS can be used to precisely simulate XY models and reveal the complex BKT phase transition.
Antiferromagnetic XY model on the triangular lattice–Antiferromagnetic interactions can induce frustration in XY models on triangular lattices [52]. Spin pairs of nearest neighbors tend to align antiparallel due to the antiferromagnetic coupling. However, spins at three vertices within each triangular unit cannot simultaneously align antiparallel to each other. It is termed geometrical frustration and leads to novel phases and orders [35]. For instance, some researches claim that this geometrical frustration in antiferromagnetic triangular lattice induces an additional second-order phase transition, and the critical temperature is slightly higher than the BKT transition temperature [53, 52]. Nevertheless, many results support the existence of only one BKT phase transition [54, 55]. To experimentally validate this, we construct the antiferromagnetic XY model on a triangular lattice by FPS, and employ Metropolis-Hastings algorithm to investigate the frustration and phase transitions.
Figure 3: Simulating the antiferromagnetic XY model on a triangular lattice with $N=12\times 12$ spins. (a) The triangular lattice (upper panel), the interacting neighbors on the SLM (lower left), and the Fourier mask for NN interactions with periodic boundary conditions (lower right). (b) The specific heat ($C_{V}$) and helicity modulus ($Y$) as functions of temperature. $C_{V}(T)$ peaks at $T_{c}=0.48J$ indicating a second-order phase transition, while $Y(T)$ intersects $\tilde{Y}(T)=\frac{2}{\pi}T$ at $T_{\text{BKT}}\approx 0.43J$, denoting a BKT transition with $T_{\text{BKT}}<T_{c}$.
The triangular lattice is obtained on a square lattice by properly deforming and assigning interactions between each spin and its six neighbors. The corresponding deformed Fourier mask is shown in FIG. 3(a) (see [34] for details). We engineer the Fourier mask to include the NN interactions and the NNN interactions along the main diagonal of the square lattice. To distinguish the two different phase transitions, we use the specific heat $C_{V}=\frac{\beta^{2}}{N}\left(\langle\mathcal{H}^{2}\rangle-\langle\mathcal{H}\rangle^{2}\right)$ to characterize the second-order phase transition [52], while use the helicity modulus to characterize the BKT phase transition. As shown in FIG. 3(b), the specific heat has a peak at $T_{c}=0.48J$ and the intersection of $Y(T)$ and $\tilde{Y}(T)$ is at the $T_{\text{BKT}}\approx 0.43J$. Our results reveal two distinct phase transitions for the antiferromagnetic XY model on a triangular lattice, in which the second-order phase transition occurs at a slightly higher temperature than the BKT phase transition, consistent with high-accuracy numerical simulations [56]. Meanwhile, the geometrical frustration of the ground states is demonstrated by the angle between any two spin pairs, which is $2\pi/3$ within each triangular unit (see [34] for one of the ground states generated by FPS). Therefore, FPS is a reliable tool for studying complex phase-transition behavior involving frustrations.
Antiferromagnetic $J_{1}$-$J_{2}$ XY model on the honeycomb lattice–To demonstrate the ability of the FPS in lattice engineering, we further study the $J_{1}$-$J_{2}$ model on the honeycomb lattice. To construct a honeycomb lattice with the square lattice pixels of the SLM, we need to stretch the lattice, shift its sites and introduce vacancies, as shown in FIG. 4(a). The NN and NNN spin interactions $J_{1}$ and $J_{2}$ are realized by Fourier masks and the ratio $J_{2}/J_{1}$ is tunable from zero to infinity [36] by changing the superposition ratio of the two corresponding Fourier masks [16]. The competition between the NN and NNN interactions leads to geometrical frustrations and highly degenerate ground states at the zero temperature. These can be characterized in reciprocal space by the distribution of spiral wave vectors within the first Brillouin zone (1BZ) [36].
Figure 4: Searching the ground state of antiferromagnetic $J_{1}$-$J_{2}$ XY model on the honeycomb lattice with $N=96$ spins. (a) The lattice structure of the honeycomb lattice (upper left), the spin arrangement on the SLM (upper right), and the Fourier masks for NN interactions ($J_{1}$, lower left) and NNN interactions ($J_{2}$, lower right) with periodic boundary conditions. In building a honeycomb lattice from a square lattice, we set the light intensity in the dotted box areas in the SLM to be zero. (b) The ground states in reciprocal space with different $J_{2}/J_{1}$ ratios. The black dashed lines mark the boundary of the 1BZ, while the white solid lines indicate the theoretically predicted positions of the peaks [36].
We experimentally found ground states with different $J_{2}/J_{1}$ ratios, as shown in FIG. 4(b). When $J_{2}=0$, only NN interaction is involved. In the ground state the nearest neighbor spin pairs are antiparallel. We obtain two sublattices with opposite spin orientations. Hence, the peak of the spiral wave vector is located at the center of the 1BZ ($\bm{\mathit{\Gamma}}$). It remains the case within the range $0\leq J_{2}/J_{1}\leq 1/6$ (FIG. 4(b1) and (b2)), indicating the Néel antiferromagnetic phase. When $1/6<J_{2}/J_{1}<1/2$, multiple peaks appear on a closed contour centered at the $\bm{\mathit{\Gamma}}$ point. The solid white lines in FIG. 4(b3) show the theoretical predictions of all the degenerate ground states. In the experiment the system will randomly relax to some of the ground states, corresponding to the separated peaks on these curves. Specifically, the peaks for $J_{2}/J_{1}=1/2$ are along white lines passing through the $\bm{M}$ points (FIG. 4(b4)). When we further increase the NNN interaction, $J_{2}/J_{1}>1/2$, the peaks appear on the closed contours around the $\bm{K}$ and $\bm{K}^{\prime}$ points (FIG. 4(b5)). When $J_{2}/J_{1}\rightarrow\infty$, the orientations of NN spin pairs in each sublattice differ by an angle $2\pi/3$, and the peaks in the 1BZ are shifted to the $\bm{K}$ and $\bm{K}^{\prime}$ points, which is a $2\pi/3$ ordered phase (FIG. 4(b6)). These experimental results are consistent with theoretical predictions [36], which demonstrates the ability of FPS for exploring spin systems with complex lattice geometries and spin interactions.
In conclusion, we present a programmable Fourier photonic spin simulator to solve the XY models with various lattice geometries and complex spin interactions. Such an FPS harnesses SLM to encode spins and the Fourier mask to obtain lattice geometries and spin interactions. This approach enables a lattice-size-independent Hamiltonian calculation for the XY models. Using FPS, we experimentally demonstrated the BKT phase transition and spin frustration in ferromagnetic square lattices, antiferromagnetic triangular lattices, and antiferromagnetic $J_{1}$-$J_{2}$ honeycomb lattices. In these models, we precisely determined the critical transition temperatures, directly observed the geometrically frustrated states, and identified the degenerate ground-state configurations. These findings validate the universality of the FPS for exploring various phases and critical behaviors across the XY models, and can be extended to quantum spin systems [57, 58]. It also has the potential to solve continuous optimization problems, accelerate statistical learning algorithms, and establish large-scale optical neural networks.
Acknowledgment–This work was supported by the National Natural Science Foundation of China (Grant No. 12325412, 12088101, 12434020), National Key Research and Development Program of China (Grant No. 2024YFA1408900), Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0303200), and “Pioneer” and “Leading Goose” R&D Program of Zhejiang (Grant No. 2025C01028) and the Fundamental Research Funds for the Central Universities.
Data availability–Digital data associated with this work are available from Ref. [34].
Y.S. and W.F. contributed equally to this work
Contact author: weiru_fan@zju.edu.cn
Contact author: dwwang@zju.edu.cn
§Contact author: hqlin@zju.edu.cn
Appendix A End Matter
XY spin encoding and Fourier mask–The SLM and camera are placed on the front and back focal planes of a lens such that the lens performs an optical Fourier transform for the modulated field. The distribution of the intensity on the camera can be described as
$I(\bm{u})=\left|\mathcal{F}\left\{\sum_{j=1}^{N}A_{j}e^{i\varphi_{j}}\operatorname{rect}\left(\frac{\bm{x}-\bm{x}_{j}}{W}\right)\right\}\right|^{2}$ (2)
where $\mathcal{F}$ means the Fourier transform, $A_{j}$ and $\varphi_{j}$ are the amplitude and phase of modulated light for the $j$-th spin, $\bm{x}=(x,~y)$ and $\bm{u}=(u,v)$ are the coordinates on the SLM and its Fourier plane of the lens, $\bm{x}_{j}$ is the coordinate of the pixel center representing the $j$-th spin, $W$ is the length of the square pixels allocated to each spin on the SLM, and $\mathrm{rect}(\bm{x})$ is the rectangular function which takes 1 if $|x|\leq 1/2$ and $|y|\leq 1/2$, otherwise 0. The phase $\varphi_{j}=\theta_{j}$ and the amplitude $A_{j}$ is proportional to the disorder strength $\xi_{j}$ in Eq. (1) (see [34] for more details).
The Hamiltonian of the XY model is obtained by making integration of $I(\bm{u})I_{M}(\bm{u})$, where $I_{M}(\bm{u})$ is a Fourier mask [16] designed for specific spin interactions. According to the convolution theorem [59], the integral of light intensity $I(\bm{u})$ multiplied by a modulation function $I_{M}(\bm{u})$ is equivalent to performing the convolution for all spins with the kernel $\mathcal{F}^{-1}\left\{I_{M}(\bm{u})\right\}$ [34]. When $\mathcal{F}^{-1}\left\{I_{M}(\bm{u})\right\}$ takes the form of summation over several delta functions, representing the interaction between a spin to others at discrete lattice sites, the corresponding Fourier mask is
$I_{M}(\bm{u})=\mathcal{F}\left\{\sum_{\alpha,\bm{r}^{(\alpha)}}J_{\alpha}\delta\left(\bm{x}-\bm{r}^{(\alpha)}\right)\right\}~,$ (3)
where $\bm{r}^{(\alpha)}$ is the distance between interacting spin sites. $J_{\alpha}$ controls the ratio of various Fourier masks by weighted sum and the summation over $\alpha$ indicates that arbitrary range interactions can be achieved by a Fourier mask. Therefore, the Hamiltonian of the XY model can be directly obtained in experiment by a correspondingly designed Fourier mask, which accelerates the sampling of spin configurations (see derivation in [34]). In our experiment, the Fourier mask is linearly renormalized to values between 0 and 1, which allows experimental implementation with an intensity attenuator such as a metasurface [60].
Computational complexity of the FPS–Due to geometrical frustration, the energy landscape of the XY model is generally rough such that it takes a long time to obtain a sufficient number of samples for large-scale lattices. FPS directly obtains the Hamiltonian of spin configurations from the measured light intensity and updates it using MCMC algorithms, which significantly speeds up the sampling process. By incorporating the Fourier-mask technique into the FPS, we can employ a single Fourier mask to realize arbitrary range spin interactions. Consequently, the speed of calculating the Hamiltonian is nearly the same for any range of spin interactions, and the computational complexity is $\mathcal{O}(N)$, while most of the current local updating algorithms scale as $\mathcal{O}(N\mathrm{log}N)$ [37] (see [34] for detailed information). This feature enables efficient calculation of large-scale XY models with complex long-range interactions [38], and can be further extended to simulate statistical models that can be mapped to XY models, including superfluids [49], superconductors [43], and Josephson junctions arrays [61]. We also provide some results on long-range interactions featuring power-law decay in square lattices (see [34]).
The time consumption of FPS includes the response time of the SLM, the acquisition time of the camera, and the computation time of Hadamard multiplication and summation between $I$ and $I_{M}$. In our implementation of the FPS, the SLM is replaced by a digital micromirror device, which has a refresh rate of 20 kHz, significantly increasing the modulation speed of the FPS [62]. However, this step also introduces a large optical aberration caused by the industrial-grade chip of digital light processing, resulting in significant energy deviations. To address this issue, we develop an iterative retrieval approach to precompensate the overall aberration of FPS (see [34] for the algorithm details and the results of the aberration compensation), enabling to precisely obtain the Hamiltonian of the XY models. For further acceleration, a metasurface [60] and photodiode can be used to replace the camera to reduce the overall time consumption. We can further speed up the sampling procedure by using state-of-the-art machine learning algorithms to suggest optimized spin-flip policies [63], which can be used to solve optimization problems in real time [64].